Given: x(x-2)=(x+4)(x+6)
Prove: x=-2
Write steps and justify property.
Please and thank you very much

Answer by stanbon(75887)

Given: x(x-2)=(x+4)(x+6)
Prove: x=-2
Write steps and justify property.
----------------------------------------
Substitute for "x":
-2(-2-2) = (-2+4)(-2+6)
Evaluate each parentheses.
-2(-4) = 2(4)
Multiply
8 = 8
Finished
=====================
Cheers,
Stan H.